if a population with size of 4 has a variance of 6.8 what is the variance of the sampling distributionof its means? the sampling distribution has a sample size of 3 and all possible samples are drawn without replacement



1
Expert's answer
2022-03-20T14:52:28-0400

"\sigma_x^2=\frac{\sigma^2}{n}
 
 =\frac{6.8}{3} 
 =2.67"